OpenStudy (thefluffmuncher):

You are the onwer of a bird store, and you use them to organize your cages. If you keep 1 bird per cage, there will be one bird outside of the cage, therefore it will be missing cages. If you keep two birds per cage, there will be one empty cage. How many birds and cages do you have? Separate the answer using commas. E.g.: 'X, Y' (X birds, Y cages)

OpenStudy (thefluffmuncher):

@Destinymasha

OpenStudy (thefluffmuncher):

@ganeshie8

OpenStudy (thefluffmuncher):

@Nnesha

OpenStudy (thefluffmuncher):

how do i solve this?

OpenStudy (thefluffmuncher):

@sammixboo

Directrix (directrix):

x = number of birds y = number of cages If you keep 1 bird per cage, there will be one bird outside of the cage means: y*1 +1 = x If you keep two birds per cage, there will be one empty cage. means: y*2 - 1 = x

Directrix (directrix):

y*1 +1 = x y*2 - 1 = x ----------- Solve these two equations simultaneously

Directrix (directrix):

@thefluffmuncher